# Mechanically Interlocked Polymers in Dilute Solution under Shear and Extensional Flows: A Brownian Dynamics Study

### v1

@@ III Results
Figure 3: Polyrotaxane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), and (c) tumbling frequency ($\omega\tau_{R}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length.
III.1 Polyrotaxanes: Steady Uniaxial Extensional Flow
At the lower end of Wi range, the behavior of polyrotaxanes diverges significantly from linear polymers. Fig. 2a shows adding rings to linear polymers increases $\langle X_{bb}\rangle/L$. This can be explained by the repulsive interactions between ring beads and backbone beads. These forces prevent backbone collapse, maintaining an extended conformation even in weak flows. Comparatively, linear polymers adopt more collapsed configurations in weak flows because of the lack of rings preventing collapse. At higher Wi values, the difference becomes negligible as both linear and polyrotaxane polymers reach fully extended backbone configurations.
The first normal stress difference $(N_{1})$ is lower in polyrotaxanes than linear polymers across the range of simulated Wi values, as shown in Fig. 2b. molecules with higher number of rings show lower $(N_{1})$ values. This can be related to the rings sliding over the backbone, leading to stress dissipation, resulting in less stress buildup in the flow direction. This shows polyrotaxanes have more pronounced viscous and weaker elastic behavior compared with linear polymers without mechanical links. The mobile rings decrease the stress response of polyrotaxanes compared to linear counterparts.
@@ III.1 Polyrotaxanes: Steady Uniaxial Extensional Flow
As shown in Fig. 2c, at low Wi, polyrotaxanes show higher extensional viscosity $(\eta_{e})$ compared to linear polymers, with higher number of rings threaded in the backbone resulting in higher $\eta_{e}$. The higher $\eta_{e}$ can be attributed to the stiffening effect of the rings on the backbone, causing the chains to be in a more extended configuration than a chain without rings. As a result, the polyrotaxanes have higher resistance to further elongation. At higher Wi values, the $\eta_{e}$ difference between polyrotaxanes and linear polymers becomes less noticeable because all chains reach a relatively extended configuration and the rings do not play a significant role in the backbone conformation and resistance to extension.
III.2 Polyrotaxanes: Steady Shear Flow
@@
Polymers in shear flow undergo continuous stretch-tumble cycles De Gennes (1974); Smith et al. (1999); Teixeira et al. (2005) and the fractional extension values reported in this work are averaged over those cycles. According to Fig. 3a, fractional extension in flow direction ($\langle X_{bb}\rangle/L$) in polyrotaxanes is higher than the linear polymers with the same number of backbone beads across the complete Wi range tested. Greater number of rings result in higher extension and more deviation from linear polymers. Overall, the repulsive ring-backbone interactions cause the molecules to exhibit extended configurations, which results in comparatively higher extension in flow direction.
A similar pattern is observed in $\langle Y_{bb}\rangle/L$ and $\langle Z_{bb}\rangle/L$, as shown in Fig. B.1a and Fig. B.1b. The presence of rings in the polyrotaxanes increase the polymer’s extent in both the gradient ($y$) and cross-flow ($z$) directions, leading to greater fractional extension values compared to linear chains without rings. This occurs because the rings effectively increase the overall spatial span of the polymer. According to Fig. B.3c, the orientation angle ($\theta$) in polyrotaxanes follows a trend similar to linear polymers, showing that the threaded rings have minimal effect on polymer orientation.
Polyrotaxanes exhibit higher shear viscosity $(\eta)$ than their linear counterparts, with a larger number of threaded rings leading to a further rise in viscosity, as shown in Fig. 3b. The increase in viscosity of polymer solutions with increasing molecular weight is a well-established relationship in polymer science Krigbaum and Flory (1953). Thereby, it is not surprising to observe adding rings to a linear chain increase the viscosity as the molecular weight is increased. In addition, rings limit the mobility of the chain, increasing the stress needed to deform the polymer from an extended configuration to induce tumbling. The power-law exponent $\beta$ is lower in polyrotaxanes, where higher number of rings have lower $\beta$ values. For example, a linear polymer with 80 beads has $\beta=-0.380$ whereas threading 4 or 16 rings to it decreases the $\beta$ value to -0.281 and -0.226, respectively, showing that polyrotaxanes have weaker shear thinning behavior which is caused by restricted segmental mobility under shear deformation caused by the rings.
@@
In weak flow ($\mathrm{Wi}\leq 1$), shear forces are not sufficient to induce significant tumbling in either polyrotaxanes or linear polymers, as shown in Fig. 3c. Consequently, both polymers have flat tumbling frequency profiles. As the flow strength increases ($101<\mathrm{Wi}<$ 10), polyrotaxanes start to show enhanced tumbling compared to the linear counterparts. At $\mathrm{Wi}\geq 10$, the threaded rings amplify the tumbling of the polyrotaxanes, causing significantly higher tumbling frequency compared with linear polymers. At $\mathrm{Wi}=800$, polyrotaxanes with 40 and 80 backbone beads (with 8 and 16 threaded rings, respectively) exhibit tumbling frequencies 38% and 35% higher than their linear polymer counterparts. Furthermore, within polyrotaxanes, threading more rings on a backbone increases the tumbling frequency. For instance, at $\mathrm{Wi}=1000$, a polyrotaxane with 80 beads in the backbone and 16 threaded rings tumbles 14% more frequently than one with only 4 rings. The enhanced tumbling in polyrotaxanes can be explained by the threaded rings increasing the polymer’s hydrodynamic cross-section along the gradient direction. The expanded profile exposes a larger surface area to shear forces of the solvent, leading to greater torque and consequently, more frequent tumbling events.
@@
Figure 5: Daisy chain properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), and (c) tumbling frequency ($\omega\tau_{R}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate higher number of segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads.
III.3 Daisy Chains: Steady Uniaxial Extensional Flow
From Fig. 4a, at low Wi in extensional flow, daisy chains and linear polymers have similar fractional extension in flow direction $\langle X_{bb}\rangle/L$. In stronger flows, daisy chains start to diverge from linear chains and show less extension. For example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) extends 55% less than a comparable linear polymer ($M_{w}=120$). The difference is less significant when the daisy chain contains longer linear segments. For example, the extension of daisy chains with 2 segments increases when beads per segment is increased from 8 to 16 and then to 32 (cold to hot colored curves), approaching the linear polymer behavior. Comparing daisy chains with the same length of linear segment and varying number of segments, shows that increasing the number of segments decreases the fractional extension, causing the chain to behave less like a linear polymer. This is due to higher fraction of the beads being in ring form, which is less extensible than a linear segment. Overall, shorter linear segments or more linear segments in a daisy chain results in more deviation from a linear polymer, resulting in lower fractional extension.
@@ III.3 Daisy Chains: Steady Uniaxial Extensional Flow
As Fig. 4b shows, at low Wi, first normal stress difference ($N_{1}$) values are close in linear and daisy chains. At higher Wi values, daisy chains have lower $N_{1}$ values compared with linear polymers with the same molecular weight. This can be explained by stress dissipation through the mobile mechanical interlocks in the daisy chains. Also, comparing a daisy chain with 8 beads per segment and 4 segments to a chain with 32 beads per segment and 2 segments, which have comparable molecular weights of 74 and 78 respectively, shows that having shorter or more sections leads to lower $N_{1}$ values (at $\mathrm{Wi}=10$, the former yields $N_{1}=577$, whereas the latter reaches a higher value of $N_{1}=1548$).
@@
Extensional viscosity ($\eta_{e}$) follows a similar trend as $N_{1}$, as shown in Fig; 4c. Daisy chains exhibit lower extensional viscosity ($\eta_{e}$) than linear polymers of equivalent molecular weight. For example, at $\mathrm{Wi}=10$, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 67% lower $\eta_{e}$ compared to a linear polymer with 80 beads. The difference is more pronounced in daisy chains with shorter or more segments. A daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has 10 times lower $\eta_{e}$ than a linear polymer with 120 beads. As an interesting example, at $\mathrm{Wi}=20$, a daisy chain with 16 beads per segment and 2 segments ($M_{w}=46$) has 55% lower extensional viscosity than a linear polymer with $M_{w}=40$, despite having 10% higher molecular weight. The same trend is evident at lower Wi values, although to a lesser extent. Overall, mechanical bonds in daisy chains lower extension, stress and $\eta_{e}$.
III.4 Daisy Chains: Steady Shear Flow
@@
As shown in Fig. 5a, fractional extension in the flow direction has different regimes based on Wi number. At $\mathrm{Wi}<0.8$, $\langle X_{bb}\rangle/L$ remains nearly constant in linear polymers and daisy chains. At higher Wi, the fractional extension increases and then decreases around $\mathrm{Wi}\approx 10^{2}$. While the conformational change is seen in both linear and daisy chains, the variation is much less significant in daisy chains with shorter or more segments. For example, a daisy chain with 8 beads per segment and 10 segments shows a nearly constant fractional extension across the whole Wi range with minimal variation, while decreasing the number of segments increases the variation. Also, in daisy chains with the same number of segments, longer segments result in behavior similar to a linear polymer. A terminal increasing regime is seen at very high Wi in linear polymers. However, this behavior could not be confirmed in daisy chains because with our current model, the chains disentangle under very strong flow conditions. We plan to address this shortcoming in future work.
According to Figs. B.2a and B.2b, fractional extensions in the gradient and cross directions show a similar constant regime at low Wi followed by a decreasing regime at higher Wi. Similar to flow direction, $\langle Y_{bb}\rangle/L$ and $\langle Z_{bb}\rangle/L$ show less variation in daisy chains with shorter or more segments. $\langle Z_{bb}\rangle/L$ in daisy chains exhibit an increasing terminal regime corresponding to the decrease in $\langle X_{bb}\rangle/L$, but we could not confirm this in all daisy chains due to model limitations. As shown by the orientation angle in Fig. B.2c, daisy chains exhibit little orientation in weak flows and reach their peak orientation angle at $\mathrm{Wi}\approx 7$, compared to $\mathrm{Wi}\approx 3$ for linear polymers, indicating that a stronger flow is needed to align daisy chains. This can be attributed to rearrangements at the mechanical bonds. At high $\mathrm{Wi}$, linear polymers reach a small angle of $\approx 2^{\circ}$, while daisy chains of equal molecular weight maintain a higher angle of $\approx 4.5^{\circ}$, reflecting differences in aspect ratio discussed in the tumbling dynamics section.
@@
Fig. 5b shows that the daisy chains are shear thinning polymers similar to linear polymers, with shear viscosity decreasing with Wi. However, the power-law exponent is lower in daisy chains with shorter or more segments, deviating from linear polymers. For example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has $\beta=-0.225$ while the corresponding linear polymer ($M_{w}=120$) has $\beta=-0.390$, which shows the less pronounced shear-thinning character of the daisy chains. This daisy chain has significantly lower viscosity as well, highlighting another major deviation from the linear polymers. For instance, at $\mathrm{Wi}=20$, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 42% lower viscosity than the corresponding linear polymer ($M_{w}=80$). It is worth noting that due to the difference in shear-thinning degree between linear and daisy chain polymers, the viscosity difference is more pronounced in weaker flows and diminishes in stronger flows. As an example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has 520% lower viscosity than the comparable linear polymer ($M_{w}=120$) at $\mathrm{Wi}=0.1$ but the difference decreases at $\mathrm{Wi}=600$ to a less significant value of  27%.
@@
Based on Fig. 5c, at low Wi, the flow is not strong enough to cause significant tumbling in either the daisy chains or linear polymers. At higher Wi, daisy chains show significantly more frequent tumbling in shear flow. At $\mathrm{Wi}=200$, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) shows 87% higher $\omega\tau_{R}$ than a linear polymer with 120 beads. Similarly, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 60% higher $\omega\tau_{R}$ than the linear counterpart with 80 beads. In shear flow, the greater extent of polymers in the gradient direction causes larger shear forces from the flow to be applied to the chains, resulting in more frequent tumbles. The rings in the daisy chains increase the gradient-flow extent. This effect is less present in configurations with fewer or longer segments, making them more linear-like in terms of aspect ratio, showing less tumbling. On the other hand, the role of the rings in the aspect ratio is more pronounced in daisy chains with more or shorter segments, causing the chains to experience larger shear forces from the flow, resulting in higher tumbling frequency. In addition to aspect ratio, the mechanical interlocks make the polymer structure more flexible, further enhancing tumbling.
@@
Figure 7: Poylcatenane properties in shear flow as a function of Weissenberg number (Wi): (a, b) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (c, d) shear viscosity ($\eta$), (e, f) tumbling frequency ($\omega\tau_{R}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$). Darker shades of green indicate increasing beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$.
III.5 Polycatenanes: Steady Uniaxial Extensional Flow
Fig. 6a shows the fractional extension of polycatenanes in flow direction. At low Wi, $\langle X_{bb}\rangle/L$ is higher in polycatenanes than linear polymers with the same number of beads. This stems from steric repulsions between interlocked rings, which enforce elongated conformations, compared with the equilibrium coiled state of linear polymers. However, at medium to high Wi, linear polymers become more extended than polycatenanes because repulsion between and within catenane rings hinders further elongation. In contrast, linear polymers uncoil and extend readily. Decreasing the number of rings increases $\langle X_{bb}\rangle/L$ because fewer rings minimize overlap, allowing the chain to approach the maximum extension of a linear chain with $N_{bb}/2$ beads. Interestingly, variation in $\langle X_{bb}\rangle/L$ is less in polycatenanes that have fewer rings. For instance, a polycatenane with 8 beads per ring and 5 rings extends only 2 times beyond equilibrium, while a polycatenane chain with the same ring size and 65 rings stretches over 3 times. This is because more rings give the chain more degrees of freedom. This can be better understood by assuming each ring a repeat unit. More repeat units in a chain makes it more flexible. In the limiting case, if we continue increasing the number of rings indefinitely, the chain approaches a linear polymer Brereton (2001).

### v2

@@ III Results
Figure 3: Polyrotaxane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.
III.1 Polyrotaxanes: Extensional Flow
At the lower end of Wi range, the behavior of polyrotaxanes diverges significantly from linear polymers. Fig. 2a shows that adding rings to linear polymers increases the fractional extension in the flow direction. This low Wi effect is largely equilibrium in nature, as demonstrated in SI Tables A.1 and A.2, which give the equilibrium radius of gyration $R_{g}$, asphericity Theodorou and Suter (1985), and prolateness Vatin et al. (2024) of linear polymers and polyrotaxanes. Shorter polyrotaxanes with a high density of mechanical bonds are more extended (higher $R_{g}$), less spherical (higher asphericity), and more rod-like (higher prolateness) than linear polymers with a similar backbone length. This is in agreement with experimental works on PEG-based polyrotaxanes where more rings elongate and rigidify the backbone Qiu et al. (2020). In general, both rings and backbones are rigidified after threading as rotaxanes, even when the starting components are flexible because the intermolecular and steric interactions reduce the conformational space and increase barriers Fadler and Flood (2022). The differences in extension become less pronounced for longer backbones or fewer mechanical bonds, which is also reflected in Fig. 2a. The low Wi differences, then, can be understood as a difference in equilibrium shape that will align with weak flows. At higher Wi values, differences due to interlocked rings become negligible as both linear and polyrotaxane polymers reach fully extended backbone configurations. Examples of linear polymers and polyrotaxanes in these steady-state stretched configurations are shown in SI Fig. C.1.
The first normal stress difference $(N_{1})$ is lower in polyrotaxanes than linear polymers across the range of simulated Wi values, as shown in Fig. 2b. Polymers with more rings exhibit lower $N_{1}$ values compared with linear polymers, demonstrating that mechanical bonds lead to weaker elastic behavior. This weaker elastic behavior is notably similar to bottlebrush polymers Dutta and Sing (2024), with increasing bristle length playing a similar role to backbone ring density in polyrotaxanes.
@@ III.1 Polyrotaxanes: Extensional Flow
As shown in Fig. 2c, at low Wi, polyrotaxanes show greater extensional viscosity $(\eta_{e})$ compared to linear polymers at the same backbone length, with higher number of rings resulting in higher $\eta_{e}$. The higher $\eta_{e}$ can be attributed to the more extended configuration of chains with threaded rings than chains without rings. As a result, polyrotaxanes have higher resistance to further elongation. At higher Wi values, the $\eta_{e}$ difference between polyrotaxanes and linear polymers lessens because all chains reach a relatively extended configuration. As shown in SI Fig. D.1, in this regime, the rings are pulled to the ends of the backbone and therefore, do not play a significant role in the backbone conformation and resistance to extension.
III.2 Polyrotaxanes: Shear Flow
@@
Polymers in shear flow undergo continuous stretch-tumble cycles De Gennes (1974); Smith et al. (1999); Teixeira et al. (2005) and the fractional extension values reported in this work are averaged over those cycles. Example configurations of a polyrotaxane (and other MIP types) during these cycles are shown in SI Fig. C.1. According to Fig. 3a, fractional extension in flow direction ($\langle X_{bb}\rangle/L$) in polyrotaxanes is higher than the linear polymers with the same number of backbone beads across the Wi range tested. A greater number of rings result in higher extension and more deviation from linear polymers. Similar to low Wi extensional flow, the ring-backbone interactions increase rigidity and cause the molecules to exhibit more elongated configurations on average, resulting in comparatively higher extension in flow direction. In real rotaxanes, this increased rigidity can be caused by hydrogen bonding between the ring and backbone, $\pi$-$\pi$ stacking between rings, metal coordination, steric interactions between rings, or electrostatic repulsions Fadler and Flood (2022).
Polyrotaxanes exhibit higher shear viscosity $(\eta)$ than their linear counterparts, with larger number of threaded rings leading to further rise in viscosity, as shown in Fig. 3b. The size of a polymer in solution has a direct relationship with the viscosity of the solution Krigbaum and Flory (1953); Hiemenz and Lodge (2007); Rubinstein and Colby (2003). As evident from SI Tables A.1 and A.2, the radius of gyration of polyrotaxanes is higher than linear polymers with the same backbone length, with additional threaded rings leading to larger polymer size. For example, a 120 bead linear polymer has a radius of gyration of 15.86, compared to a polyrotaxane with 24 rings on a 120 bead backbone, which has a radius of gyration of 23.80.
Similar to linear polymers, polyrotaxanes show shear thinning, although to a lesser degree, as evident from the power-law exponent $\beta$ values shown in SI Table E.1. Additionally, polyrotaxanes with more rings have lower $\beta$ values. For example, a linear polymer with 80 beads has $\beta=-0.380$ (in agreement with the previously reported value for this model Dutta and Sing (2024)) whereas threading 4 or 16 rings on it decreases the $\beta$ value to 0.290-0.290 and 0.225-0.225, respectively, showing that more rings further suppress shear thinning. The weaker shear thinning is due to the polyrotaxanes’ decreased alignment with the flow, demonstrated by the average orientation angle in SI Fig. B.1c. For $\mathrm{Wi}>1$, increasing the number of rings results in less alignment with the flow on average (an angle $\theta$ with greater deviation from $0^{\circ}$) because the rings cannot fully align with the flow due to repulsive interactions within the rings and between the rings and the backbone. This misalignment leads to less relaxation and weaker shear thinning Teixeira et al. (2005); Ryder and Yeomans (2006).
@@
In weak flow ($\mathrm{Wi}\leq 1$), shear forces are not sufficient to induce significant tumbling in either polyrotaxanes or linear polymers, as shown in Fig. 3c. Consequently, both polymers have flat tumbling frequency profiles. As the flow strength increases ($1<\mathrm{Wi}<10$), polyrotaxanes start to show enhanced tumbling compared to the linear counterparts. At $\mathrm{Wi}\geq 10$, the polyrotaxanes show significantly higher tumbling frequency compared with linear polymers. At $\mathrm{Wi}=800$, polyrotaxanes with 40 and 80 backbone beads (with 8 and 16 threaded rings, respectively) exhibit tumbling frequencies 38% and 35% higher than their linear polymer counterparts. Furthermore, within polyrotaxanes, threading more rings on a backbone increases the tumbling frequency. For instance, at $\mathrm{Wi}=10^{3}$, a polyrotaxane with 80 beads in the backbone and 16 threaded rings tumbles 14% more frequently than one with only 4 rings. The enhanced tumbling in polyrotaxanes can be explained by the threaded rings increasing the polymer’s hydrodynamic cross-section along the gradient direction, as shown is SI Fig. B.1a. The expanded profile exposes a larger surface area to shear forces of the solvent, leading to more frequent tumbling events. The polyrotaxanes also have a greater extent in the cross-flow direction, demonstrated in SI Fig. B.1b.
Fig. 3d shows the first normal stress coefficient ($\psi_{1}$) as a function of Wi. At a comparable backbone length, polyrotaxanes have higher $\psi_{1}$ than linear polymers, with more rings further increasing the value. As shown in SI Table E.2, the power-law exponent ($\beta$) in polyrotaxanes is only slightly higher than linear polymers. For example, for a polyrotaxane with 120 beads in the backbone and 24 threaded rings, $\beta=-1.24$ while for a linear polymer with $N=120$, we report $\beta=-1.32$. The latter value is in close agreement with the previous reported value of -1.30 for this model Dutta and Sing (2024).
@@
Figure 5: Daisy chain properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate higher number of segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.
III.3 Daisy Chains: Extensional Flow
From Fig. 4a, at low Wi in extensional flow, daisy chains are more extended than linear polymers at similar $M_{w}$ with higher density of mechanical bonds leading to greater fractional extension. As an example, at Wi=0.1, a daisy chain with 8 beads per segment and 4 segments ($M_{w}=74$) is 72% more extended than a comparable linear polymer ($M_{w}=80$) whereas a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$), which has lower density of mechanical bonds, is only 22% more extended. This is reflected in the shape parameters of these two polymers in equilibrium, where the chain with more short segments is more prolate (0.74 vs. 0.49) and has higher asphericity (0.70 vs. 0.61), demonstrated in SI Tables A.1 and A.3.
This behavior is similar to backbone rigidifying effect of the rings as seen in the polyrotaxanes. In chains with longer segments, this effect diminishes and the chains become more coiled like a linear polymer. Interestingly, shorter linear segments lead to greater extension at low Wi, but the trend reverses at high Wi where daisy chains with longer linear segments are more extended. In strong flow regime where the chains are far from equilibrium and more extended, the repulsions between rings becomes important and a stronger flow should be needed to stretch daisy chains with more segments. Making segments longer diminishes this effect and the chains become more like a linear polymer.
@@ III.3 Daisy Chains: Extensional Flow
As Fig. 4b shows, at low Wi, first normal stress difference ($N_{1}$) values are close in linear and daisy chains. At higher Wi values, daisy chains have lower $N_{1}$ values compared with linear polymers with the same molecular weight. Also, comparing a daisy chain with 8 beads per segment and 4 segments to a chain with 32 beads per segment and 2 segments, which have comparable molecular weights of 74 and 78 respectively, shows that higher density of mechanical bonds leads to lower $N_{1}$ values. At $\mathrm{Wi}=10$, the former yields $N_{1}=577$, whereas the latter reaches a higher value of $N_{1}=1548$.
@@
Extensional viscosity ($\eta_{e}$) follows a similar trend as $N_{1}$, as shown in Fig. 4c. Daisy chains exhibit lower $\eta_{e}$ than linear polymers of equivalent molecular weight. For example, at $\mathrm{Wi}=10$, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 67% lower $\eta_{e}$ compared to a linear polymer with 80 beads. Higher mechanical bond density increases the difference. A daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has an extensional viscosity 10 times lower than a linear polymer with 120 beads. The same trend is evident at lower Wi values, but to a lesser extent. Overall, mechanical bonds reduce stress and $\eta_{e}$, but increase extension at low Wi while decreasing extension at high Wi.
III.4 Daisy Chains: Shear Flow
@@
As shown in Fig. 5a, extension in the flow direction ($\langle X_{bb}\rangle/L$) has different regimes depending on flow strength. Previous studies on semi-flexible linear polymers show a similar shear response where polymers stretch at low shear rates and collapse at high shear rates Lyulin et al. (1999). At $\mathrm{Wi}<0.8$, $\langle X_{bb}\rangle/L$ remains nearly constant in linear polymers and daisy chains, with daisy chains exhibiting greater extension than linear polymers of similar molecular weight. For example, at $\mathrm{Wi}=10^{-1}$, a daisy chain of 6 linear segments with 8 beads each ($M_{w}=118$) has a fractional extension of 0.16 compared to a linear polymer of $M_{w}=120$ with an extension of 0.09. At $10^{0}<\mathrm{Wi}<10^{2}$, $\langle X_{bb}\rangle/L$ increases and then decreases around $\mathrm{Wi}\approx 10^{2}$. While the conformational change is seen in both linear and daisy chains, the variation is less significant in daisy chains with more mechanical bonds. For example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has a fractional extension between 0.16 and 0.27 across the Wi range compared to a linear polymer of $M_{w}=120$, which has extensions between 0.08 and 0.029 across the same Wi range. Decreasing the number of segments or increasing the length of the linear segments increases the variation in extension. Also, in daisy chains with the same number of segments, longer segments result in behavior similar to a linear polymer. A terminal regime with increasing extension is seen at very high Wi in linear polymers. However, this behavior could not be confirmed in daisy chains because with our current model, the chains disentangle under very strong flow conditions due to bond crossing. We plan to address this shortcoming in future work.
@@
Fig. 5b shows that the shear thinning degree is lower in daisy chains compared to linear polymers and the power-law exponent $\beta$ (SI Table E.1) is lower in daisy chains with higher mechanical bond density. For example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has $\beta=-0.225$ while the corresponding linear polymer ($M_{w}=120$) has $\beta=-0.390$, which shows the less pronounced shear-thinning character of the daisy chains. Shear-thinning behavior becomes more linear-like as the daisy chains’ linear segment length increases. Similar to polyrotaxanes, this decreased shear-thinning in daisy chains is caused by their misalignment with the flow direction. SI Fig. B.2c shows that the orientation angle deviates from the flow direction more for daisy chains with shorter linear segments or fewer segments. At a comparable molecular weight, daisy chains also have lower zero-shear viscosity due to their smaller equilibrium size, as shown in SI Table A.3, highlighting another significant deviation from linear polymers. For example, at $\mathrm{Wi}=20$, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 42% lower viscosity than the corresponding linear polymer ($M_{w}=80$). It is worth noting that because of the difference in shear-thinning degree between linear and daisy chain polymers, the viscosity difference is more pronounced in weaker flows and diminishes in stronger flows. As an example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has 520% lower viscosity than the comparable linear polymer ($M_{w}=120$) at $\mathrm{Wi}=0.1$ but the difference decreases at $\mathrm{Wi}=600$ to a less significant value of  27%.
@@
Based on Fig. 5c, at low Wi, the flow is not strong enough to cause significant tumbling in either the daisy chains or linear polymers. At higher Wi (>7>7), daisy chains show significantly more frequent tumbling in shear flow. At $\mathrm{Wi}=200$, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) shows 87% higher $\omega\tau_{R}$ than a linear polymer with 120 beads. Similarly, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 60% higher $\omega\tau_{R}$ than the linear counterpart with 80 beads. In shear flow, the greater extent of polymers in the gradient direction (shown in SI Fig. B.2a) causes larger shear forces from the flow to be applied to the chains, resulting in more frequent tumbles. The rings in the daisy chains increase the gradient-flow extent. This effect is less present in configurations with lower density of mechanical bonds, making them more linear-like in terms of aspect ratio, showing less tumbling. On the other hand, the role of the rings in the aspect ratio is more pronounced in daisy chains with higher mechanical bond density, causing the chains to experience larger shear forces from the flow, resulting in higher tumbling frequency.
According to Fig. 5d, daisy chains show lower $\psi_{1}$ compared with linear polymers at the same molecular weight, with higher density of mechanical bonds further decreasing $\psi_{1}$ value. As shown in SI Table E.2, the power-law exponent ($\beta$) for daisy chains can be lower or higher than linear polymers with comparable molecular weight depending on the density of mechanical bonds. For example, for a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$), $\beta=-1.349$ which is lower than the value of -1.239 for a linear polymer with 80 beads. On the other hand, a daisy chain with 8 beads per segment and 4 segments ($M_{w}=74$) has $\beta=-1.105$ due to higher density of mechanical bonds.
@@
Figure 7: Poylcatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$). Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.
Figure 8: Poylcatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Darker shades of green indicate increasing beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.
III.5 Polycatenanes: Extensional Flow